### v1

@@ 4.4 Telescopic extension
Let $\ell\in\{0,...,L\}$, $L\in\mathbb{N}^{+}$ represent a series of accuracy levels associated with both the inflated spaces $\mathcal{V}_{\varepsilon}^{k,\ell}\subseteq\mathcal{V}$ and the deflated spaces $\mathcal{W}_{\varepsilon}^{k,\ell}\subseteq\mathcal{V}$. These follow a nested inclusion relationship, $\mathcal{V}_{\varepsilon}^{k,0}\subseteq...\subseteq\mathcal{V}_{\varepsilon}^{k,L-1}\subseteq\mathcal{V}$, and $\mathcal{W}_{\varepsilon}^{k,0}\subseteq...\subseteq\mathcal{W}_{\varepsilon}^{k,L-1}\subseteq\mathcal{V}_{\varepsilon}^{k,L-1}$, with $\mathcal{V}_{\varepsilon}^{k,L}=\mathcal{W}_{\varepsilon}^{k,L}=\mathcal{V}$, for all $k\in\{0,...,N_{T}\}$. This hierarchy of spaces corresponds to a sequence of models $\Psi_{\varepsilon}^{k,\ell}(\,\cdot\,,t_{k+1}-t_{k}):\mathcal{V}^{k,\ell}\rightarrow\mathcal{V}^{k,\ell}$ and of paired ensembles $(E_{k|k}^{P,0},E_{k|k}^{C,0}\equiv\emptyset)$, $(E_{k|k}^{P,\ell},E_{k|k}^{C,\ell})$, defined as follows
$E_{k|k}^{P,\ell}$$\coloneqq\{\omega_{k|k}^{P,n,\ell}\in\mathcal{V}_{\varepsilon}^{k,\ell}\}_{n=1}^{M_{\ell}},\quad$$\forall\ell\in\{0,...,L\}$
$E_{k|k}^{C,\ell}$$\coloneqq\{\omega_{k|k}^{C,n,\ell}=\Pi_{\mathcal{V}_{\varepsilon}^{k,\ell-1}}\omega_{k|k}^{P,n,\ell}\in\mathcal{V}_{\varepsilon}^{k,\ell-1}\}_{n=1}^{M_{\ell}},\quad$$\forall\ell\in\{1,...,L\}.$
Here $M_{\ell}$ denotes the size of the ensembles at the $\ell$th accuracy level, for any $\ell\in\{0,...,L\}$. The principal ensemble at the 00th level coincides with the ancillary ensemble of the two-level case.
@@
$\mathcal{V}_{\varepsilon}^{k+1,\ell}=\text{span}\{\text{POD}(\Xi_{\Delta,k}^{\text{HF}},\varepsilon_{k}^{\Diamond,\ell}),\mathcal{W}_{\varepsilon}^{k,\ell}\},\qquad\forall\ell\in\{0,...,L-1\},$
@@
$\mathcal{W}_{\varepsilon}^{k+1,\ell}=\text{POD}(\Xi_{\Delta,k}^{\text{LF}},\varepsilon_{k}^{\Diamond,\ell}),\qquad\forall\ell\in\{0,...,L-1\},$
@@
$\varepsilon^{\text{ML},\ell}_{k}=\dfrac{2\varepsilon_{r}^{\ell}}{M_{0}}\sum_{n=1}^{M_{0}}\|\omega_{k|k}^{P,n,0}-\overline{\omega}_{k|k}^{P,0}\|_{\mathcal{V}}^{2}+\sum_{s=1}^{L}\dfrac{2\varepsilon_{r}^{\ell}}{M_{s}}\sum_{n=1}^{M_{s}}\Big{(}\|\omega_{k|k}^{P,n,s}-\overline{\omega}_{k|k}^{P,s}\|_{\mathcal{V}}^{2}$
@@
$\varepsilon_{k}^{\text{MF},\ell}=\dfrac{2\varepsilon_{r}^{\ell}}{4^{L}M_{0}}\sum_{n=1}^{M_{0}}\|\omega_{k|k}^{P,n,0}-\overline{\omega}_{k|k}^{P,0}\|_{\mathcal{V}}^{2}+\sum_{s=1}^{L}\dfrac{2\varepsilon_{r}^{\ell}}{4^{L-s}M_{s}}\sum_{n=1}^{M_{s}}\Big{(}\|\omega_{k|k}^{P,n,s}-\overline{\omega}_{k|k}^{P,s}\|_{\mathcal{V}}^{2}$
@@ 5 Numerical experiments on quasi-geostrophic equations
The weak formulation of problem (19) reads: find $\omega\in C^{1}([0,T];L^{2}(\Omega))\cap C^{0}([0,T];$ $H^{1}_{0}(\Omega))$ and $\psi\in C^{0}([0,T];H^{1}_{0}(\Omega))$ such that, for all $\eta\in H^{1}_{0}(\Omega)$ and $\xi\in H^{1}_{0}(\Omega)$,
@@
At each iteration, a reduced basis space is constructed employing POD, detailed in Section 3.1, on a training set of vorticity fields. This leads to an approximation space $\mathcal{V}_{\varepsilon}^{k}\coloneqq\text{span}\{\phi_{1},...,\phi_{N_{\varepsilon}^{k}}\}$ for the vorticity field, and to an approximation space $\mathcal{U}_{\varepsilon}^{k}=\text{span}\{\zeta_{1},...,\zeta_{N_{\varepsilon}^{k}}\}$ for the stream function. These basis functions of the two spaces satisfy the relationship
@@
The semi-discretization gives rise to a set of nonlinear ordinary differential equations that are solved using the implicit midpoint rule. The time integration involves dividing the time span $[0,T]$ into subintervals of length $dt=0.1$. In our context, we set $T=5000$, and we designate the solution over the time range $[1000,2000]$ as the ‘true’ state trajectory. The solution at time, $t_{0}=1000$, is shown in Figure 1. To characterize the attractor of the dynamics, we consider the system’s trajectory over the long term horizon $[4000,5000]$. The collection of snapshots computed over this interval allows the construction of a reduced space denoted as $\mathcal{V}_{\varepsilon}\subset P^{1}_{h}$ and enables the sampling of states from an approximation of the invariant measure.
@@ 5.1 Numerical results
$\text{err}(t_{k}|\mathbf{d}_{r})=\dfrac{\|\overline{\omega}^{P}_{k|r}-\omega^{\star}(t_{k})\|_{L^{2}(\Omega)}}{\|\omega^{\star}(t_{k})\|_{L^{2}(\Omega)}},\quad r\in\{k,k+1\},\quad\forall k=\{0,...,999\}.$
@@
The error calculation is repeated for four trajectory reconstructions employing independent noise realizations. The resulting confidence intervals enable the distinction between the stochastic fluctuations and the mean behavior of the filters. As further performance indicators, we monitor how the size of the reduced model changes over the assimilation time and the computational time required for the algorithm to execute each assimilation step.
@@
As a first experiment, we assess the performance of the standard ensemble Kalman against those of the multi-level and multi-fidelity ensemble Kalman filters. For the latter, we consider two levels of accuracy and we employ different approaches to construct surrogate models. Notably, we introduce a reference multi-level filter and a reference multi-fidelity filter that employ the high-fidelity model in place of the surrogate model. Although the resulting filter is prohibitively expensive for practical applications, it does not introduce approximation errors and thus it serves as a reference for the other multi-accuracy algorithms. The considered multi-level and multi-fidelity algorithms encompass the inflation-deflation aRB-ML-EnKF and aRB-MF-EnKF methods, and a filter that makes use of memory-less reduced order modeling, detailed in donoghue22 and already mentioned in Remark 2.
@@
All filters employ ensembles initialized with states sampled from a Gaussian distribution with inverse Laplacian covariance, as detailed in Section 2.1. The standard EnKF ensemble has $N_{P}=48$ high-fidelity states, whereas the multi-accuracy filters have $N_{P}=48$ principal states, $N_{C}=48$ control states, and $N_{A}=1200$ ancillary states. The relative accuracy for the surrogate models is set to $\varepsilon_{r}=10^{-3}$.
@@
Figure 2: This plot compares different data assimilation algorithms. In the left column are reported the results of filters employing multi-level updates, while in the right column those using of multi-fidelity updates. The top row reports the relative errors versus assimilation time, the center row displays how the reduced basis size changes over the assimilation time, and the bottom row shows the computational time per assimilation step versus assimilation time.
@@
In Figure 2, we present the results of this experiment. The error plots show good performances of the inflation-deflation aRB-ML-EnKF and aRB-MF-EnKF, whose error closely aligns to the error of the reference algorithm, both in the multi-level and in the multi-fidelity case, and it is 2 to 2.5 orders of magnitude smaller than the errors obtained with the standard EnKF and with the memory-less algorithm. The evolution of the size of the surrogate model reveals that the inflation-deflation approach ensures larger and more accurate reduced spaces compared to the memory-less algorithm. This comes at the price of a higher computational cost in the first few assimilation windows that, however, becomes asymptotically comparable to the one of the memory-less algorithm, and that can be reduced by choosing a better prior distribution, as shown in Figure 3. Finally, we remark that, for the same computational cost, the multi-level update produces slightly more accurate reconstructions than the multi-fidelity one, although the behavior of the two algorithms does not differ significantly.
@@
Figure 3: This plot compares different sampling priors. In the left column are reported the results of filters employing multi-level updates, while in the right column those using of multi-fidelity updates. The sub-plots in the top row illustrate relative errors versus assimilation time, those in the center row display the relationship between reduced basis size and assimilation time, and those at bottom the computational times per assimilation step versus assimilation time.
The results of the experiment, summarized in Figure 3, demonstrate the benefits of integrating additional information about the system dynamics into the data assimilation process. Using the invariant measure of dynamics as a prior leads to significantly faster convergence for both algorithms. However, it also poses a challenge for the adaptive reduced order modeling scheme. Specifically, the error plot reveals over the last assimilation windows a one order of magnitude difference in relative errors between the inflation-deflation aRB-ML-EnKF/aRB-MF-EnKF method, and the reference algorithm. To better understand this discrepancy, we can examine the plots depicting the surrogate model size (central row of Figure 3). There, we observe a rapid contraction of the approximation space during the first assimilation steps, suggesting a degradation in the model’s approximation properties. This information loss is mitigated by the inflation-deflation approach, which extracts additional information from the system dynamics, thereby expanding the size of the approximation space. Nevertheless, this compensatory mechanism does not appear sufficient to close the gap that arises in the early assimilation steps.
@@
Figure 4: This plot compares different tolerances for the surrogate model. In the left column are reported the results of filters employing multi-level updates, while in the right column those using of multi-fidelity updates. The sub-plots in the top row illustrate relative errors versus assimilation time, those in the center row display the relationship between reduced basis size and assimilation time, and those at bottom the computational times per assimilation step versus assimilation time.
@@
In the last experiment, we compare the state reconstructions obtained employing the inflation-deflation ML-EnKF with ensembles of varying sizes. Specifically, we explore the impact of the principal and of the ancillary ensemble size on the reconstruction error, on the surrogate model size, and on the computational cost. For this analysis, we first consider a scenario where the principal and control ensemble sizes are fixed, $N_{P}=N_{C}=48$, while the ancillary ensemble size $N_{A}$ varies in $\{600,1200,2400,4800\}$. Subsequently, we explore the scenario where the ancillary ensemble size is fixed, $N_{A}=1200$, and the principal and control ensemble sizes vary, $N_{P}=N_{C}\in\{12,48,192,768\}$. In both cases, we maintain a relative tolerance of $\varepsilon_{r}=10^{-3}$ and we employ initial ensembles sampled from a Gaussian prior with inverse Laplacian covariance.
@@
Figure 5: This plot compares the results of the inflation-deflation ML-EnKF algorithm obtained by employing ensembles of different sizes. In the left column, varying values of the ancillary ensemble size are considered, while in the right column, varying values of the principal and control size are considered. The sub-plots in the top row illustrate relative errors versus assimilation time, those in the center row display the relationship between reduced basis size and assimilation time, and those at bottom the computational times per assimilation step versus assimilation time.

### v2

@@ 4.4 Telescopic extension
Let $\ell\in\{0,...,L\}$, $L\in\mathbb{N}^{+}$ represent a series of accuracy levels associated with both the inflated spaces $\mathcal{V}_{\varepsilon}^{k,\ell}\subseteq\mathcal{V}$ and the deflated spaces $\mathcal{W}_{\varepsilon}^{k,\ell}\subseteq\mathcal{V}$. These follow a nested inclusion relationship, $\mathcal{V}_{\varepsilon}^{k,0}\subseteq...\subseteq\mathcal{V}_{\varepsilon}^{k,L-1}\subseteq\mathcal{V}$, and $\mathcal{W}_{\varepsilon}^{k,0}\subseteq...\subseteq\mathcal{W}_{\varepsilon}^{k,L-1}\subseteq\mathcal{V}_{\varepsilon}^{k,L-1}$, with $\mathcal{V}_{\varepsilon}^{k,L}=\mathcal{W}_{\varepsilon}^{k,L}=\mathcal{V}$, for all $k\in\{0,...,N_{T}\}$. This hierarchy of spaces corresponds to a sequence of models ${\color[rgb]{0,0,0}\Psi_{\varepsilon}^{k,\ell}(\,\cdot\,,t_{k},t_{k+1})}:\mathcal{V}^{k,\ell}\rightarrow\mathcal{V}^{k,\ell}$ and of paired ensembles $(E_{k|k}^{P,0},E_{k|k}^{C,0}\equiv\emptyset)$, $(E_{k|k}^{P,\ell},E_{k|k}^{C,\ell})$, defined as follows
$E_{k|k}^{P,\ell}$ :=$:=\{\omega_{k|k}^{P,n,\ell}\in\mathcal{V}_{\varepsilon}^{k,\ell}\}_{n=1}^{M_{\ell}},\quad$$\forall\ell\in\{0,...,L\}$
$E_{k|k}^{C,\ell}$ :=$:=\{\omega_{k|k}^{C,n,\ell}=\Pi_{\mathcal{V}_{\varepsilon}^{k,\ell-1}}\omega_{k|k}^{P,n,\ell}\in\mathcal{V}_{\varepsilon}^{k,\ell-1}\}_{n=1}^{M_{\ell}},\quad$$\forall\ell\in\{1,...,L\}.$
Here $M_{\ell}$ denotes the size of the ensembles at the $\ell$th accuracy level, for any $\ell\in\{0,...,L\}$. The principal ensemble at the 00th level coincides with the ancillary ensemble of the two-level case. The size of each ensemble should be determined considering the error and costs of the surrogate models. Ideas on how to manage this operation and maximize the efficiency of multi-level and multi-fidelity Monte Carlo methods can be found in jakeman2020 ; oljava2018 ; peherstorfer2016 ; schaden2020 .
@@
$\mathcal{V}_{\varepsilon}^{k+1,\ell}=\text{span}\{\text{POD}(\Xi_{\Delta,k}^{\text{HF}},\varepsilon_{k}^{\Diamond,\ell}/2),\mathcal{W}_{\varepsilon}^{k,\ell}\},\qquad\forall\ell\in\{0,...,L-1\},$
@@
$\mathcal{W}_{\varepsilon}^{k+1,\ell}=\text{POD}(\Xi_{\Delta,k}^{\text{LF}},\varepsilon_{k}^{\Diamond,\ell}/2),\qquad\forall\ell\in\{0,...,L-1\},$
@@
$\varepsilon^{\text{ML},\ell}_{k}=\dfrac{2\varepsilon_{r}^{\ell}}{M_{0}-1}\sum_{n=1}^{M_{0}}\|\omega_{k|k}^{P,n,0}-\overline{\omega}_{k|k}^{P,0}\|_{\mathcal{V}}^{2}+\sum_{s=1}^{L}\dfrac{2\varepsilon_{r}^{\ell}}{M_{s}-1}\sum_{n=1}^{M_{s}}\Big{(}\|\omega_{k|k}^{P,n,s}-\overline{\omega}_{k|k}^{P,s}\|_{\mathcal{V}}^{2}$
@@
$\varepsilon_{k}^{\text{MF},\ell}=\dfrac{2\varepsilon_{r}^{\ell}/4^{L}}{M_{0}-1}\sum_{n=1}^{M_{0}}\|\omega_{k|k}^{P,n,0}-\overline{\omega}_{k|k}^{P,0}\|_{\mathcal{V}}^{2}+\sum_{s=1}^{L}\dfrac{2\varepsilon_{r}^{\ell}/4^{L-s}}{M_{s}-1}\sum_{n=1}^{M_{s}}\Big{(}\|\omega_{k|k}^{P,n,s}-\overline{\omega}_{k|k}^{P,s}\|_{\mathcal{V}}^{2}$
@@ 5 Numerical experiments on quasi-geostrophic equations
The weak formulation of problem (21) reads: find $\omega\in C^{1}([0,T];L^{2}(\Omega))\cap C^{0}([0,T];$ $H^{1}_{0}(\Omega))$ and $\psi\in C^{0}([0,T];H^{1}_{0}(\Omega))$ such that for all $\eta,\xi\in H_{0}^{1}(\Omega)$
@@
At each iteration, a reduced basis space is constructed employing POD, detailed in Subsection 3.1, on a training set of vorticity fields. This leads to an approximation space $\mathcal{V}_{\varepsilon}^{k}:=\text{span}\{\phi_{1},...,\phi_{N_{\varepsilon}^{k}}\}$ for the vorticity field, and to an approximation space $\mathcal{U}_{\varepsilon}^{k}=\text{span}\{\zeta_{1},...,\zeta_{N_{\varepsilon}^{k}}\}$ for the stream function. These basis functions of the two spaces satisfy the relationship
@@
The semi-discretization gives rise to a set of nonlinear ordinary differential equations that are solved using the implicit midpoint rule. The time integration involves dividing the time span $[0,T]$ into subintervals of length $dt=0.1$. In our context, we set $T=5000$, and we designate the solution over the time range $[1000,2000]$ as the ‘true’ state trajectory. The solution at time, $t_{0}=1000$, is shown in Figure 1. To characterize the attractor of the dynamics, we consider the system’s trajectory over the long term horizon $[4000,5000]$. The collection of snapshots computed over this interval allows the construction of global reduced basis models. Additionally, randomized selection of these snapshots serves as a sampling strategy, generating states that are distributed according to the invariant measure of the dynamics.
@@ 5.1 Numerical results
$\text{err}(t_{k}|\mathbf{d}_{r})=\dfrac{\|\overline{\omega}^{P}_{k|r}-\omega^{\star}(t_{k})\|_{L^{2}(\Omega)}}{\|\omega^{\star}(t_{k})\|_{L^{2}(\Omega)}},\quad r\in\{k,k+1\},\quad\forall k=\{0,...,1000\}.$
@@
The error calculation is repeated for four trajectory reconstructions employing independent noise realizations. The resulting confidence intervals enable the distinction between the stochastic fluctuations and the mean behavior of the filters. As further performance indicators, we monitor how the size of the reduced model changes over the assimilation time and the wall clock time required for the algorithm to execute each data assimilation step, i.e., the time required to complete one prediction and one update step. For the post-processing, we also consider the cumulative computational cost, i.e., the sum of the individual costs associated with all the data assimilation steps up to a given assimilation time, and the long-term computational cost, i.e. the average cost of the last ten data assimilation steps before a given assimilation time.
@@
As a first experiment, we assess the performance of the standard ensemble Kalman filter against those of the multi-level and multi-fidelity ensemble Kalman filters under identical long-term computational cost conditions. For the multi-level and multi-fidelity filters, we consider two levels of accuracy and we employ different approaches to construct surrogate models. Notably, we introduce a reference multi-level filter and a reference multi-fidelity filter that employ the high-fidelity model in place of the surrogate model. Although the resulting filter is prohibitively expensive for practical applications, it does not introduce approximation errors and thus it serves as a reference for the other multi-accuracy algorithms. The considered multi-level and multi-fidelity algorithms encompass the inflation-deflation aRB-ML-EnKF and aRB-MF-EnKF methods, and a filter that makes use of memory-less reduced order modeling, detailed in donoghue22 and already mentioned in Remark 2.
@@
All filters employ ensembles initialized with states sampled from a Gaussian distribution with inverse Laplacian covariance, as detailed in Subsection 2.1. The standard EnKF ensemble has $N_{P}=256$ high-fidelity states, whereas the multi-accuracy filters have $N_{P}=48$ principal states, $N_{C}=48$ control states, and $N_{A}=1200$ ancillary states. The relative accuracy for the surrogate models is set to $\varepsilon_{r}=10^{-3}$.
@@
Figure 2: This plot compares different data assimilation algorithms. The left column shows the results of filters employing multi-level updates, while the right column shows those using multi-fidelity updates. The top row reports the relative errors versus assimilation time, the center row displays how the reduced basis size changes over the assimilation time, and the bottom row shows the wall clock time per single data assimilation step versus assimilation time.
@@
In Figure 2, we present the results of this experiment. The error plots show good performances of the inflation-deflation aRB-ML-EnKF and aRB-MF-EnKF, whose error closely aligns to the error of the reference algorithm, both in the multi-level and in the multi-fidelity case, and it is 1 to 1.5 orders of magnitude smaller than the errors obtained with the standard EnKF and with the memory-less algorithm. The evolution of the size of the surrogate model reveals that the inflation-deflation approach ensures larger and more accurate reduced spaces compared to the memory-less algorithm. However, this comes with a higher computational cost in the initial assimilation windows, which can be mitigated by selecting a better prior distribution, as shown in Figure 3. Finally, we remark that, for the same computational cost, the multi-level update produces slightly more accurate reconstructions than the multi-fidelity one, although the behavior of the two algorithms does not differ significantly.
@@
Figure 3: This plot compares different sampling priors. The left column shows the results of filters employing multi-level updates, while the right column shows those using multi-fidelity updates. The sub-plots in the top row illustrate relative errors versus assimilation time, those in the center row display the relationship between reduced basis size and assimilation time, and those at bottom the wall clock time per single data assimilation step versus assimilation time.
The results of the experiment, summarized in Figure 3, demonstrate the benefits of integrating additional information about the system dynamics into the data assimilation process. Using the invariant measure of dynamics as a prior leads to significantly faster convergence for both algorithms. However, it also poses a challenge for the adaptive reduced order modeling scheme. Specifically, the error plot reveals over the last assimilation windows a one order of magnitude difference in relative errors between the inflation-deflation aRB-ML-EnKF/aRB-MF-EnKF method, and the reference algorithm. To better understand this discrepancy, we can examine the plots depicting the surrogate model size (central row of Figure 3). There, we observe a rapid contraction of the approximation space during the first data assimilation steps, suggesting a degradation in the model’s approximation properties. This information loss is mitigated by the inflation-deflation approach, which extracts additional information from the system dynamics, thereby expanding the size of the approximation space. Nevertheless, this compensatory mechanism does not appear sufficient to close the gap that arises in the early data assimilation steps.
@@
Figure 4: This plot compares different tolerances for the surrogate model. The left column shows the results of filters employing multi-level updates, while the right column shows those using multi-fidelity updates. The sub-plots in the top row illustrate relative errors versus assimilation time, those in the center row display the relationship between reduced basis size and assimilation time, and those at bottom the wall clock time per single data assimilation step versus assimilation time.
@@
Figure 5: This plot compares different data assimilation algorithms: the standard EnKF, the two-level aRB-ML-EnKF, and three RB-EnKF with different accuracy levels and ensemble sizes. In the left column are reported the results of filters employing the inverse Laplacian prior, while in the right column those using the invariant measure prior. The top row reports the relative errors versus assimilation time, the center row displays how the reduced basis size changes over the assimilation time, and the bottom row shows the wall clock time per single data assimilation step versus assimilation time.
In a fourth experiment, we compare the state reconstruction obtained with the standard EnKF and with the aRB-ML-EnKF to those obtained using a RB-EnKF that employs global reduced basis models of different accuracy. These models are constructed offline, leveraging the same set of full order solutions used to obtain samples from the invariant measure of the dynamics. For the aRB-ML-EnKF, we consider two levels of accuracy, while for the other filters, we use a single level of accuracy. We use initial ensembles sampled from a Gaussian distribution with inverse Laplacian covariance and from the same invariant measure used for the offline construction of the RB-EnKF surrogate models. The size of all ensembles is chosen to ensure the same long-term computational cost per data assimilation step. Specifically, for the standard EnKF we consider ensembles of $N_{P}=256$ high-fidelity states, whereas for the multi-level filter, we use a principal ensemble of size $N_{P}=48$, an ancillary ensemble of size $N_{A}=1200$, and set the relative accuracy to $\varepsilon_{r}=10^{-3}$. For the global RB-EnKF, we employ three surrogate models with absolute accuracy $\varepsilon_{a}\in\{10^{-7},10^{-5},10^{-3}\}$ corresponding to low-fidelity ensembles of size $N_{P}\in\{400,800,1600\}$, respectively.
In Figure 5, we present the results of the experiment. Using low-fidelity models in place of high-fidelity models significantly reduces the convergence time of the EnKF, especially when the surrogate models have lower accuracy. This effect can be explained by the fact that the global reduced basis models retain information about the system dynamics, allowing the filter to operate within a lower dimensional approximation space. However, for both priors used, this faster convergence has a limit on achievable accuracy, with less accurate surrogate models leading to higher reconstruction errors. This error stagnation aligns with the accuracy of the approximation spaces considered. When comparing these RB-EnKF filters to our proposed adaptive approach, the adaptive method appears to be more accurate in the long run. Specifically, in the case of an invariant measure prior, where all methods benefit from knowledge of the system’s long-term behavior, both the standard EnKF and the aRB-ML-EnKF outperform the three global RB-EnKF methods in terms of achieved reconstruction error, with the aRB-ML-EnKF being the most accurate.
Figure 6: This plot compares different data assimilation algorithms: the standard EnKF, the two-level aRB-ML-EnKF, and three two-level RB-ML-EnKF with offline-trained low-fidelity models. The left column shows the results of filters employing the inverse Laplacian prior, while in the right column those using the invariant measure prior. The top row reports the relative errors versus assimilation time, the center row displays how the reduced basis size changes over the assimilation time, and the bottom row shows the wall clock time per single data assimilation step versus assimilation time.
In the next experiment, we compare the state reconstruction obtained with the standard EnKF and with the aRB-ML-EnKF to those obtained from three multi-level ensemble Kalman filters employing the same global low-fidelity models used in the previous experiment. We consider two levels of accuracy for all the filters except the standard EnKF, and we use the same inverse Laplacian and measure invariant priors used before. The size of the ensembles is again chosen to ensure the same long-term computational cost for the filters. For the standard EnKF and for the aRB-ML-EnKF, we consider the same testing conditions as from the previous experiment, while for the global RB-ML-EnKF, we employ principal ensembles of full order states of size $N_{P}=48$, and ancillary ensembles of size $N_{A}\in\{330,660,1320\}$, respectively associated to the global reduced basis model with accuracy $\varepsilon_{a}\in\{10^{-7},10^{-5},10^{-3}\}$.
The results of this experiment are presented in Figure 6. Contrary to the previous test, the convergence of the global RB-ML-EnKF algorithms does not seem to be affected by the accuracy of the global low-fidelity models. However, the accuracy achieved with the non-adaptive multi-level filters is comparable to that of the standard EnKF at the same cost, i.e., one to two orders of magnitude less accurate than the reconstruction obtained with the proposed adaptive method.
@@
In the sixth experiment, we compare the state reconstructions obtained employing the inflation-deflation ML-EnKF with ensembles of varying sizes. Specifically, we explore the impact of the principal and of the ancillary ensemble size on the reconstruction error, on the surrogate model size, and on the computational cost. For this analysis, we first consider a scenario where the principal and control ensemble sizes are fixed, $N_{P}=N_{C}=48$, while the ancillary ensemble size $N_{A}$ varies in $\{600,1200,2400,4800\}$. Subsequently, we explore the scenario where the ancillary ensemble size is fixed, $N_{A}=1200$, and the principal and control ensemble sizes vary, $N_{P}=N_{C}\in\{12,48,192,768\}$. In both cases, we maintain a relative tolerance of $\varepsilon_{r}=10^{-3}$ and we employ initial ensembles sampled from a Gaussian prior with inverse Laplacian covariance.
@@
Figure 7: This plot compares the results of the inflation-deflation ML-EnKF algorithm obtained by employing ensembles of different sizes. In the left column, varying values of the ancillary ensemble size are considered, while in the right column, varying values of the principal and control size are considered. The sub-plots in the top row illustrate relative errors versus assimilation time, those in the center row display the relationship between reduced basis size and assimilation time, and those at bottom the wall clock time per single data assimilation step versus assimilation time.
@@
Figure 8: The plot compares the reconstruction error and computational cost of three different data assimilation algorithms. On the left: the relative error versus the cumulative clock time required for all data assimilation steps. On the right: the relative error versus the clock time of the last data assimilation step.
In the last experiment, we compare the accuracy and computational cost of the standard EnKF with those of the inflation-deflation aRB-ML-EnKF and aRB-MF-EnKF. For the standard EnKF, we consider ensembles of size $N_{P}\in\{48,300,600,1200,2400,4800\}$, while for the multi-level and multi-fidelity filters, we use two levels of accuracy, with a principal ensemble of fixed size $N_{P}=48$, and ancillary ensembles of size $N_{A}\in\{48,600,1200,2400,4800,9600\}$. The states in the ensemble of the standard EnKF and in the principal ensemble of the multi-fidelity and multi-level EnKF are forecasted using the high-fidelity model. The results, presented in Figure 8, display the reconstruction error after $1000$ data assimilation steps, compared to the cumulative cost of the algorithms up to the final time $t_{1000}$ (left plot), and the cost of a single data assimilation step at the final time (right plot). The aRB-ML-EnKF appears to be the least costly of the methods, approaching the mean-field limit faster than the standard EnKF. For example, to achieve an accuracy of $10^{-5}$ at the final time, the standard EnKF requires a cumulative computational cost that is double that of the aRB-ML-EnKF. In contrast, the aRB-MF-EnKF exhibits a more complex behavior: initially, with an ancillary ensemble of less than 2400 states, the algorithm is more accurate than the standard EnKF for the same cost. However, for larger ensemble sizes and higher cumulative costs, its performance significantly deteriorates. This behavior, rather than indicating slower convergence to the same mean-field limit, seems to align with the predictions of Remark 1: the MF-EnKF does not appear to share the same mean-field limit as the other two filters, making it, at least in this example, a less effective filter.